OpenStudy (rayaz):

Use the binomial theorem to find the exact value of \[\left( \sqrt{3}+1 \right)^{5}-\left( \sqrt{3}-1 \right)^{5}\] Hence show, without using tables or calculators, that the value of \[\left( \sqrt{3}+1 \right)^{5}\] lies between 152 and 153.

OpenStudy (trojanpoem):

Number of terms = n+1 = 5+1 = 6 (1+sqrt(3))^5 + (1- sqrt(3))^5 (1+sqrt(3))^5 = 1 - sqrt(3) * 5 + sqrt(3) * 5 * 4 / 2! - sqrt(3) * 5 * 4 * 3 /3! + sqrt(3) * 5 * 4 * 3 * 2/3! - sqrt(3) * 5 * 4 * 3 * 2 * 1/4! (1- sqrt(3))^5= 1 + sqrt(3) * 5 + sqrt(3) * 5 * 4 / 2! + sqrt(3) * 5 * 4 * 3 /3! + sqrt(3) * 5 * 4 * 3 * 2/3! + sqrt(3) * 5 * 4 * 3 * 2 * 1/4! So we will get 2 * ( the odd terms) , we could have induced that at first, as It's common in binomials. 2 ( 1 + sqrt(3)*5*4/2! + sqrt(3) * 5 * 4 * 3 * 2 * 1/(3!) ) << 6/2 = 3 terms (2 + 20sqrt(3) + 40sqrt(3) )

OpenStudy (rayaz):

Shouldn't your final answer be \[2+20\left( \sqrt{3} \right)^{2}+10\left( \sqrt{3} \right)^{4}\]

ganeshie8 (ganeshie8):

Hey!

OpenStudy (rayaz):

Hi

ganeshie8 (ganeshie8):

\[\left( \sqrt{3}+1 \right)^{5}-\left( \sqrt{3}-1 \right)^{5} \\~\\ = \sum\limits_{k=0}^5 \dbinom{5}{k}(\sqrt{3})^k (1)^{5-k} - \sum\limits_{k=0}^5 \dbinom{5}{k}(\sqrt{3})^k (-1)^{5-k}\\~\\ = \sum\limits_{k=0}^5 \dbinom{5}{k}(\sqrt{3})^k [(1)^{5-k}-(-1)^{5-k}] \\~\\ = \sum\limits_{k=0}^5 \dbinom{5}{k}(\sqrt{3})^k [1-(-1)^{5-k}] \\~\\ \] Clearly, the terms cancel out when \(k\) is odd. Therefore the sum is : \[2\left[\dbinom{5}{0}(\sqrt{3})^0+\dbinom{5}{2}(\sqrt{3})^2+\dbinom{5}{4}(\sqrt{3})^4\right]\] simplify

OpenStudy (rayaz):

Thanks for the explanation, can you help with the second part of the question. I don't know how to approach it. This is ow I've been trying to solve it \[\left( \sqrt{3}+1 \right)^{5}\ =152+\left( \sqrt{3}-1 \right)^{5}\] After this I'm stuck.

ganeshie8 (ganeshie8):

Very good!

ganeshie8 (ganeshie8):

for second part, just notice this : \(\sqrt{3}-1\) is between \(0\) and \(1\)

OpenStudy (rayaz):

Oh! Well that was clumsy of me to not notice. Thanks.

ganeshie8 (ganeshie8):

\[0\lt \sqrt{3}-1\lt 1\] \[0\lt (\sqrt{3}-1)^5\lt 1\] \[152+0\lt 152+(\sqrt{3}-1)^5\lt 152+ 1\]

ganeshie8 (ganeshie8):

There is another nice way to work this problem with out using binomial theorem

ganeshie8 (ganeshie8):

If you're interested, I can give you one or two hints :)

OpenStudy (rayaz):

Go ahead

ganeshie8 (ganeshie8):

Hint 1 : Notice that \(1\pm \sqrt{3}\) are the roots of quadratic \(x^2-2x-2=0\)

ganeshie8 (ganeshie8):

Let \(r = 1+\sqrt{3}\) and \(s=1-\sqrt{3}\) Since these are the roots of above quadratic, we have : \[s^2-2s-2=0\implies s^2 = 2s+2\tag{1}\]\[r^2-2r-2=0\implies r^2 = 2r+2\tag{2}\]

ganeshie8 (ganeshie8):

multiplying both sides by \(s^{k-2}\) and \(r^{k-2}\) gives \[s^k = 2s^{k-1}+2s^{k-2}\tag{3}\]\[ r^k = 2r^{k-1}+2r^{k-2}\tag{4}\]

ganeshie8 (ganeshie8):

Adding (3) and (4) : \[s^k+r^k = 2(s^{k-1}+r^{k-1})+2(s^{k-2}+r^{k-2})\tag{5}\]

ganeshie8 (ganeshie8):

we already know that \(s+r = 2\) plugin \(k=2\) in (5) : \[s^2+r^2 \\= 2(s^{2-1}+r^{2-1})+2(s^{2-2}+r^{2-2})\\=2(s+r)+2(1+1)\\=2(2)+2(2)\\=8\]

ganeshie8 (ganeshie8):

plugin \(k=3\) in (5) : \[s^3+r^3 \\= 2(s^{3-1}+r^{3-1})+2(s^{3-2}+r^{3-2})\\=2(s^2+r^2)+2(s+r)\\=2(8)+2(2)\\=20\]

ganeshie8 (ganeshie8):

plugin \(k=4\) in (5) : \[s^4+r^4 \\= 2(s^{4-1}+r^{4-1})+2(s^{4-2}+r^{4-2})\\=2(s^3+r^3)+2(s^2+r^2)\\=2(20)+2(8)\\=56\]

ganeshie8 (ganeshie8):

plugin \(k=5\) in (5) : \[s^5+r^5 \\= 2(s^{5-1}+r^{5-1})+2(s^{5-2}+r^{5-2})\\=2(s^4+r^4)+2(s^3+r^3)\\=2(56)+2(20)\\=152\]

ganeshie8 (ganeshie8):

As you can see we can find sum of power of roots recursively using above trick

OpenStudy (rayaz):

I'm gonna try this the next time I see a similar question. Once again, thanks a bunch.

ganeshie8 (ganeshie8):

np :) do feel free to tag me in all your future questions !

OpenStudy (rayaz):

Ok